# On the Hilbert Space of Dyons

### v1

@@ 1 Introduction
Our interest in this problem was motivated by a series of papers by Csaki et. al. csaba1 ; csaba2 , following earlier work of Zwanziger zwan ; zwan0 . They argued that the usual description of multi-particle Fock spaces as the tensor product of single particle Hilbert spaces needs some refinement when the particles in question are dyons. This refinement comes in the form of what these authors called “pairwise helicity”, an extra phase that dyonic particles pick up under a rotation $R\in SU(2)$. Specifically, the rotation acts on elements of the Fock space as
@@
with ${\bf p}_{i}^{\prime}=R{\bf p}_{i}$. There is a great deal of redundancy in the phase $\varphi$ because there’s no canonical way to compare the phases of states that correspond to different rays in the Hilbert space. Indeed, the sole invariant information arises when the $\mathbf{p}_{i}$ are colinear and we consider a rotation with $\mathbf{p}^{\prime}_{i}=\mathbf{p}_{i}$: then the phase $\varphi$ is physical.
@@
For $n=2$ particles, the existence of the phase follows in a reasonably straightforward way by looking at the profiles of the electric and magnetic fields and seeing how they transform under the little group that leaves the relative momentum ${\bf p}_{1}-{\bf p}_{2}$ unchanged. This is seen most clearly in the work on Wu and Yang wuyang which we review later. For $n\geq 3$ particles, however, the calculation is less direct. It was intuited by Zwanziger as an obvious generalisation of the two-particle case, and subsequently derived by studying a kind of doubled field theory in which both electric and magnetic potentials are introduced before projection onto a particular duality frame zwan0 .
The purpose of this paper is to provide a clean derivation of the pairwise helicity formula (1.1). We do this by looking not at the Fock space of relativistic field theories, but instead at non-relativistic quantum mechanics. We will show that the additional phase is a direct consequence of topology. This arises because the wavefunctions that describe dyons are not functions, but sections of a line bundle that may be topologically non-trivial. We start just with a collection of $n$ non-relativistic particles and show that the formula (1.1) naturally arises as the most general transformation of a collection of particles under rotation. In addition, we will also derive an expression for a lower bound on the relative angular momentum carried by a collection of dyons.
@@ 2.2 Topology and Dyons
Note also that the winding numbers $w_{ij}$ are duality invariant and the $U(1)$ connection over $\mathcal{M}_{n}$ should not be though of as magnetic, or electric: it too is duality invariant. But the action (2.6) includes only the Lorentz force law, and not the Coulomb repulsion which must be put in separately. That is not surprising: the Lorentz force law and Coulomb law are related only by relativistic physics and we are firmly in the non-relativistic realm. If we were to add a Coulomb term, we would break the duality symmetry and specify the individual charges $\vec{e}_{i}$ and $\vec{g}_{i}$, rather than just the combinations $w_{ij}$.
@@ 2.4 Constructing the Bundles by Pullback
$F=\frac{1}{4}\epsilon^{\alpha\beta\gamma}\sum_{i<j}\frac{\mathbf{x}_{ij}^{\gamma}}{|\mathbf{x}_{ij}|^{3}}d\mathbf{x}_{ij}^{\alpha}\wedge d\mathbf{x}_{ij}^{\beta}\ ,$ (2.32)
@@ Acknowledgements
Thanks to Jack Smith for helpful guidance towards the classification of multi-particle bundles, and to Ofri Telem for very helpful comments on a draft of this paper. Thanks also to Nick Manton and Ben Gripaios for helpful discussions. This work was supported by STFC grant ST/L000385/1, the EPSRC grant EP/V047655/1 “Chiral Gauge Theories: From Strong Coupling to the Standard Model”, and a Simons Investigator Award. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
@@ Appendix B Proof of Bundle Construction
is an isomorphism. Let us see why this is true. $\pi_{2}(\mathcal{M}_{n})$ is a free Abelian group generated by elements $\rho_{ij}$ for $i<j$, defined by moving the $i^{\text{th}}$ point around a small $S^{2}$ enclosing only the $j^{\text{th}}$ point. Conversely, $\pi_{2}(\widehat{\mathcal{M}}_{2})$ is a free Abelian group generated by elements $\sigma_{ij}$ for $i<j$, defined by moving $\mathbf{r}_{ij}$ in an $S^{2}$ about the origin in its ${\mathbb{R}}^{3}\setminus\{\mathbf{0}\}$ while holding all other coordinates fixed. Observe then that $\sigma_{ij}$ is homotopic to $\phi\circ\rho_{ij}$, and thus $\phi_{*}\rho_{ij}=\sigma_{ij}$. We have therefore established that (B.73) is indeed an isomorphism.

### v2

@@ 1 Introduction
Our interest in this problem was motivated by a series of papers by Csaki et. al. csaba1 ; csaba2 , following earlier work of Zwanziger zwan ; zwan0 . They argued that the usual description of multi-particle Fock spaces as the tensor product of single particle Hilbert spaces needs some refinement when the particles in question are dyons. This refinement comes in the form of what these authors called “pairwise helicity”, an extra phase that dyonic particles pick up under a rotation $R\in SU(2)$. Specifically, the rotation acts on elements of the Fock space, with elements labelled by the momenta ${\bf p}_{i}$, as
@@
where ${\bf p}_{i}^{\prime}=R{\bf p}_{i}$. There is a great deal of redundancy in the phase $\varphi$ because there’s no canonical way to compare the phases of states that correspond to different rays in the Hilbert space. Indeed, the sole invariant information arises when the $\mathbf{p}_{i}$ are colinear and we consider a rotation with $\mathbf{p}^{\prime}_{i}=\mathbf{p}_{i}$: then the phase $\varphi$ is physical.
@@
For $n=2$ particles, the existence of the phase follows in a reasonably straightforward way by looking at the profiles of the electric and magnetic fields and seeing how they transform under the little group that leaves the relative momentum ${\bf p}_{1}-{\bf p}_{2}$ unchanged. This is seen most clearly in the work of Wu and Yang wuyang which we review later. For $n\geq 3$ particles, however, the calculation is less direct. It was intuited by Zwanziger as an obvious generalisation of the two-particle case, and subsequently derived by studying a kind of doubled field theory in which both electric and magnetic potentials are introduced before projection onto a particular duality frame zwan0 . (See, also, brown1 ; brown2 ; yale for further commentary on these ideas.)
The purpose of this paper is to provide a clean derivation of the pairwise helicity formula (1.1). We do this by looking not at the Fock space of relativistic field theories, but instead at non-relativistic quantum mechanics. We will show that the additional phase is a direct consequence of topology. This arises because the wavefunctions that describe dyons are not functions, but sections of a line bundle that may be topologically non-trivial. We start just with a collection of $n$ non-relativistic particles and show that the formula (1.1) naturally arises as the most general transformation under a rotation. In addition, we will also derive an expression for a lower bound on the relative angular momentum carried by a collection of dyons.
@@ 2.2 Topology and Dyons
Note also that the winding numbers $w_{ij}$ are duality invariant and the $U(1)$ connection over $\mathcal{M}_{n}$ should not be thought of as magnetic, or electric: it too is duality invariant. But the action (2.6) includes only the Lorentz force law, and not the Coulomb repulsion which must be put in separately. That is not surprising: the Lorentz force law and Coulomb law are related only by relativistic physics and we are firmly in the non-relativistic realm. If we were to add a Coulomb term, we would break the duality symmetry and specify the individual charges $\vec{e}_{i}$ and $\vec{g}_{i}$, rather than just the combinations $w_{ij}$.
@@ 2.4 Constructing the Bundles by Pullback
$F=\frac{1}{4}\epsilon^{\alpha\beta\gamma}\sum_{i<j}\frac{w_{ij}\mathbf{x}_{ij}^{\gamma}}{|\mathbf{x}_{ij}|^{3}}d\mathbf{x}_{ij}^{\alpha}\wedge d\mathbf{x}_{ij}^{\beta}\ ,$ (2.32)
@@ Acknowledgements
Thanks to Jack Smith for helpful guidance towards the classification of multi-particle bundles, and to Ofri Telem for very helpful comments on a draft of this paper. Thanks also to Nick Manton and Ben Gripaios for helpful discussions, as well as Philip Boyle Smith for comments on an earlier version. This work was supported by STFC grant ST/L000385/1, the EPSRC grant EP/V047655/1 “Chiral Gauge Theories: From Strong Coupling to the Standard Model”, and a Simons Investigator Award. For the purpose of open access, the author has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.
@@ Appendix B Proof of Bundle Construction
is an isomorphism. Let us see why this is true. $\pi_{2}(\mathcal{M}_{n})$ is a free Abelian group generated by elements $\rho_{ij}$ for $i<j$, defined by moving the $i^{\text{th}}$ point around a small $S^{2}$ enclosing only the $j^{\text{th}}$ point; a proof of this fact can be found for instance in Sec II of fadell . Conversely, $\pi_{2}(\widehat{\mathcal{M}}_{2})$ is a free Abelian group generated by elements $\sigma_{ij}$ for $i<j$, defined by moving $\mathbf{r}_{ij}$ in an $S^{2}$ about the origin in its ${\mathbb{R}}^{3}\setminus\{\mathbf{0}\}$ while holding all other coordinates fixed. Observe then that $\sigma_{ij}$ is homotopic to $\phi\circ\rho_{ij}$, and thus $\phi_{*}\rho_{ij}=\sigma_{ij}$. We have therefore established that (B.73) is indeed an isomorphism.